# gamecenter,boxed

### v1

@@ 6.1. Experimental Setup
Iris (Fisher, 1988) 150 4 3
LSun (Ultsch, 2005) 400 2 3
S1 (Fränti and Virmajoki, 2006) 5000 2 15
House (Fränti and Sieranoja, 2018) 1837 3 3
Adult (Becker and Kohavi, 1996) 48842 6 3
Birch2 (Zhang et al., 1997) 25000 2 100
SynthOV 7.5K ~ 12.5K 2-20 2-20
SynthSEP 7.5K ~ 12.5K 2-20 2-20
@@
6.2. Utility Evaluation
In this section, we focus on evaluating the utility of FastLloyd compared to the state-of-the-art work of Su et al. (Su et al., 2016). We begin by describing the metrics used for this evaluation.
Metrics
Following similar literature (Su et al., 2016), we use the Normalized Intra-cluster Variance (NICV) metric to evaluate clustering quality. In short, NICV is the $k$-means objective function (1) divided by the number of points in the dataset.
$NICV(O)=\frac{1}{N}\cdot\sum_{j=1}^{k}\sum_{x_{l}\in O_{j}}||x_{l}-\mu_{j}||^{2}$
We also introduce another metric, Empty Clusters, that measures the ratio of empty clusters to $k$. The objective of $k$-means clustering is to produce exactly $k$ clusters such that the WCSS (NICV) is minimized. However, existing metrics/objectives do not capture the requirement of having exactly $k$ clusters. This becomes problematic in differentially private settings, where adding noise can isolate some centroids from the data completely, resulting in a suboptimal solution that uses less than $k$ clusters. We introduce this metric to precisely quantify this effect as follows:
Empty $\texttt{Empty Clusters}(O)=\frac{\sum_{j=1}^{k}{I(|O_{j}|=0)}}{k}$
6.2.1. Real World Datasets
To answer Q1 and Q2, we evaluate FastLloyd, CSuLloyd, and SuLloyd on a variety of datasets. Figure 3 shows the NICV against various privacy budgets ($\epsilon$) from 0.01 to 2.0 on six datasets. The shading shows the 95% confidence interval over the 20 experiments.
First, we consider Q1. We observe that FastLloyd outperforms SuLloyd on all datasets, especially for lower values of epsilon. On a larger dataset (birch2), Figure 3(c) shows a statistically significant improvement (the confidence intervals do not overlap). FastLloyd even outperforms the non-private baseline of Lloyd at $\epsilon=2$. For higher dimensional datasets, a substantial performance gap is evident; Figure 3(f) shows that even in high-privacy settings (with $\epsilon=0.01$), FastLloyd significantly outperforms the non-private baseline of Lloyd. To answer Q2, we observe that for all datasets CSuLloyd  outperforms SuLloyd. However, in all cases, FastLloyd outperforms SuLloyd (except in Figure 3(f) , where we tie).
To evaluate the Empty Clusters metric, we use the birch2 dataset with a relatively high cluster count of 100. Figure 4 shows that even the non-private Lloyd loses around $15\%$ of clusters, with SuLloyd reaching up to more than $40\%$ in high-privacy settings (with $\epsilon=0.01$). We observe that FastLloyd eliminates this problem completely for $\epsilon>0.75$ and significantly diminishes it for smaller privacy budgets.
Figure 4. Birch2 Empty Clusters
@@ 6.2.1. Real World Datasets
(c) Improvement Percentage
Figure 5. NICV Utility comparison over SynthOV dataset. Averaged over 60 runs (20 on each of three random datasets).
(a) SuLloyd
(b) FastLloyd
(c) Improvement Percentage
Figure 6. NICV Utility comparison over SynthSEP dataset. Averaged over 60 runs (20 on each of three random datasets).
6.2.2. Synthetic Datasets
To answer Q3, we follow an approach similar to Su et al. (Su et al., 2016) and evaluate FastLloyd and SuLloyd on SynthSEP and SynthOV, varying both $k$ and $d$. Figures 5 and 6 show heatmaps of NICV values for both approaches when $\epsilon=1$. To clarify the difference between the protocols, Figures 5(c) and 6(c) show the percentage of reduction in NICV of FastLloyd over that of SuLloyd. For the overlapping dataset (SynthOV), FastLloyd  outperforms SuLloyd in all cases, with the improvement being more significant in higher cluster counts and low dimensions, reducing NICV by more than 40%. The gap in performance becomes more pronounced in the separated dataset (SynthSEP), with FastLloyd  significantly outperforming SuLloyd in all cases, reducing the NICV error by up to 89% in some cases, and by more than 78% even in the highest number of clusters and dimensions. These observations highlight the better scalability and utility of FastLloyd over SuLloyd in high-dimensional and high-cluster-count scenarios. Since SuLloyd was found to outperform all other approaches for $d{}>3$ (Su et al., 2016), we claim that FastLloyd is state-of-the-art in DP $k$-means for $d{}>3$. FastLloyd also fixes the scalability issue that Su et al.’s work (Su et al., 2016) observed in high $k$.
6.3. Runtime Evaluation
@@
$N$ $k$ $d$ FastLloyd MohLloyd (Mohassel et al., 2020) Speedup FastLloyd MohLloyd (Mohassel et al., 2020) Reduction
TimeSynth 10K 2 2 $3.23\pm 0.23$ 10500 3250×3250$\times$3250 × 128 2.557e8 1,998,000×1,998,000$\times$1 , 998 , 000 ×
2 5 $8.61\pm 1.81$ 320 - -
5 2 $18.02\pm 3.59$ 34050 1890×1890$\times$1890 × 320 9.746e8 3,045,000×3,045,000$\times$3 , 045 , 000 ×
5 5 $19.83\pm 2.01$ 800 - -
TimeSynth 100K 2 2 $85.57\pm 2.42$ 105120 1228×1228$\times$1228 × 128 2.467e9 19,277,000×19,277,000$\times$19 , 277 , 000 ×
2 5 $313.47\pm 6.79$ 320 - -
5 2 $423.1\pm 6.88$ 347250 820×820$\times$820 × 320 9.535e9 29,797,000×29,797,000$\times$29 , 797 , 000 ×
5 5 $412.89\pm 5.84$ 800 - -
LSun 400 3 2 $1.81\pm 0.1$ 1481 1156×1156$\times$1156 × 192 - -
S1 5K 15 2 $24.76\pm 0.3$ 49087 1982×1982$\times$1982 × 960 - -
@@
Table 3. Overhead comparison per iteration against MohLloyd (Mohassel et al., 2020) for two clients ($20$ runs with mean and $95\%$ confidence reported for FastLloyd)
@@
To answer questions Q4 and Q5, we compare FastLloyd with the MohLloyd protocol of Mohassel et al. (Mohassel et al., 2020) on a variety of common datasets in terms of runtime and communication size (per iteration). Table 3 shows the summary of the evaluation with two clients (the setting considered in Mohassel et al.’s work (Mohassel et al., 2020)). Values for MohLloyd are taken as reported in the paper by Mohassel et al. (Mohassel et al., 2020), noting that the setup is similar to the one used in our evaluation, and the gap in performance is more than what could be accounted for by different setups. In terms of runtime, MohLloyd executes in the order of minutes, while FastLloyd executes in a fraction of a second, offering four orders of magnitude speedup. In terms of communication size, MohLloyd requires communicating gigabytes of data per iteration, while FastLloyd needs a fraction of a kilobyte, offering up to seven orders of magnitude reduction in size. In terms of communication rounds, MohLloyd requires $\Theta(\lceil\log{k{}}\rceil)$ communication rounds per iteration (Hegde et al., [n. d.]), while FastLloyd only requires one. Since MohLloyd was found to be the state-of-the-art in secure federated $k$-means (Hegde et al., [n. d.]), we conclude that we advance the state-of-the-art while offering output privacy (which would only further slow down Mohassel et al. (Mohassel et al., 2020)).
7. Extensions
7.1. Distance-Constrained $k$-means
(a) S1 NICV
(b) Birch2 NICV
(c) Birch2 Empty Clusters
Figure 7. The effect of the distance constraint (dashed lines) on the clustering quality of FastLloyd and CSuLloyd
[…]
Figure 8. An example reduction of an assignment instance ($N=5,k=3$) to a network flow graph for $\boldsymbol{\eta}^{min}$, $\boldsymbol{\eta}^{max}$ and $\eta^{dist}$.
We introduce a new extension of the constrained clustering problem that further tunes and improves the utility of DP clustering. This new problem includes a constraint on the maximum radius ($\eta_{d}$) of every cluster, such that the objective function is now subject to the constraints:
(12) $\eta_{min}\leq|O_{j}|\leq\eta_{max}\quad\text{and}\quad||x_{l}-\mu_{j}||\leq\eta_{d}\quad\forall x_{l}\in O_{j},\forall j$
In lieu of previous work, we extend the network flow formulation to solve our new constrained clustering problem; an example is shown in Figure 8. In short, we modify the capacity of edges connecting the data points to the centroids to be $1$ if the distance is less than or equal to $\eta_{d}$, and 00 otherwise, essentially removing the edge. This prevents data points from being assigned to centroids that are further than $\eta_{d}$ away, even if it means that they do not get assigned to any cluster.
Intuitively, this constraint has several benefits. First, it provides a form of resistance to outliers, as data points that are far away from any centroid will not (and in some cases, should not) be assigned to any cluster. Second, it allows for more fine-grained control over the clusters, as the maximum distance constraint can be used to ensure that the clusters are compact. Most importantly, the maximum distance constraint helps to tighten the sensitivity bound of the centroids even further.
Theorem 7.1.
If constraints (12) are satisfied, then:
(13) $\Delta^{\mu}\leq\frac{2\eta_{d}}{\eta_{min}}$
@@ 7.1. Distance-Constrained $k$-means
Since the update phase treats every dimension separately (see Section 5.3, we can consider only one dimension. Given a previous centroid $\mu_{t-1}$ (that might be randomly initialized), we need to bound the maximum change in the new centroid $\mu_{t}$, when changing the neighbouring dataset. Specifically, we bound the following
(14) $\Delta^{\mu}=\max\limits_{D\sim D^{\prime}\in\mathcal{D}}||\mu_{t}(D,\mu_{t-1})-\mu_{t}(D^{\prime},\mu_{t-1})||_{1}.$
In the unbounded model, $D\sim D^{\prime}$ means that the datasets are identical except for one data point, $x^{\prime}$, which is in $D^{\prime}$ but not in $D$.
If $x^{\prime}$ is not within $\eta_{d}$ of $\mu_{t-1}$, it will not be assigned to the new cluster, and the neighbouring centroids will be the same, completing the proof. Therefore, we only need to consider the case where $x^{\prime}$ is within $\eta_{d}$ of $\mu_{t-1}$. In this case, we have two neighbouring assignments, $O_{t}$ and $O_{t}^{\prime}=O_{t}\cup\{x^{\prime}\}$, and the corresponding centroids will be:
$\mu_{t}$ $=$ $\frac{\sum_{x_{l}\in O_{t}}x_{l}}{|O_{t}|}$
$\mu_{t}^{\prime}$ $=$ $\frac{x^{\prime}+\sum_{x_{l}\in O_{t}}x_{l}}{|O_{t}|+1}$
Now, letting $\Sigma=\sum_{x_{l}\in O_{t}}x_{l}$, and $C=|O_{t}|$ to simplify the notation, we bound the difference between the neighbouring centroids as follows:
(15) $||\mu_{t}-\mu_{t}^{\prime}||_{1}$ $=\left|\left|\frac{\Sigma}{C}-\frac{\Sigma+x^{\prime}}{C+1}\right|\right|_{1}=\left|\left|\frac{\Sigma-Cx_{l}}{C(C+1)}\right|\right|_{1}$
(16) $\leq\left|\left|\frac{\Sigma-C(\mu_{t}\pm 2\eta_{d})}{C(C+1)}\right|\right|_{1}$
(17) $=\left|\left|\frac{\Sigma-C(\frac{\Sigma}{C}\pm 2\eta_{d})}{C(C+1)}\right|\right|_{1}$
(18) $=\left|\left|\frac{\mp 2\eta_{d}C}{C(C+1)}\right|\right|_{1}\leq\frac{2\eta_{d}}{C}\leq\frac{2\eta_{d}}{\eta_{min}}$
where (18) follows from the fact that $C\geq\eta_{min}$ and (16) follows from the fact that $x^{\prime}$ is within $\eta_{d}$ of $\mu_{t-1}$, and therefore by the triangle inequality:
$||x_{l}-x^{\prime}||_{2}\leq||x_{l}-\mu_{t-1}||_{2}+||\mu_{t-1}-x^{\prime}||_{2}\leq 2\eta_{d}\quad\forall x_{l}\in O_{t}$
which implies that $x^{\prime}$ is within $2\eta_{d}$ of any point in $O_{t}$, namely their average $\mu_{t}$. Since the $\ell_{1}$ norm over one dimension ($\ell_{\infty}$) is always less than or equal to the $\ell_{2}$ norm, we have:
$||x^{\prime}-\mu_{t}||_{1}\leq||x^{\prime}-\mu_{t}||_{2}\leq 2\eta_{d}$
implying that, in this dimension, $x^{\prime}=\mu_{t}\pm 2\eta_{d}$ in the worst case. ∎
We heuristically set this parameter as follows (for 2 dimensions):
(19) $\eta_{d}=\sqrt{\frac{B^{2}}{2t}}$
The intuition behind this choice is that it allows a larger distance ($\eta_{d}\approx 0.7B$) at the beginning of the algorithm (to allow the initial centroids to align with the data) and then gradually reduces the distance, to reduce the noise and fine-tune the clusters. In particular, we propose that $\eta_{d}$ should be a function of $k$ and $d$, as the higher number of clusters implies the more compact they are expected to be, and the higher number of dimensions implies that the data points are likely to be more spread out.
A preliminary evaluation using the S1 and Birch2 datasets shows that this constraint can significantly improve the utility of the algorithm. Figure 7 shows this improvement on both the constrained sum-count method CSuLloyd and the constrained centroid method FastLloyd, where we observe a $30\sim 40\%$ reduction in NICV in small $\epsilon$. The trade-off with this maximum radius constraint is that the network flow becomes harder to satisfy. We find that for other datasets, such as Adult, the min-cost flow solver does not always converge. We leave further exploration of this constraint to future work.
7.2. Serverless Extensions
In this section, we discuss alternatives to using the service provider. A trivial alternative is that each client could locally noise their centroids before sending them to another client (or set of clients) to aggregate. The downside to this approach is that the error will compound with the number of clients and decrease utility. An alternative is for all or a subset of the clients to take the role of a computation node in an MPC protocol. In this case, the centroids would be secret shared and aggregated in MPC, then Laplace noise would need to be added similar to FastLloyd. The noise addition can be implemented by each computation node sampling noise locally and adding it to their shares. Depending on how many nodes are assumed to be honest, this will necessarily add more noise than FastLloyd to ensure privacy in the presence of one or more nodes colluding. Alternatively, the computation nodes can jointly sample the noise using the protocol of Wu et al. (Wu et al., 2016). The disadvantage of this approach is that it will add a large computational overhead which will make the runtime closer to that of Mohassel et al. (Mohassel et al., 2020).
8. Conclusion
In this work, we designed FastLloyd, a new private protocol for federated $k$-means. We have shown that FastLloyd is secure in the computational model of DP and analyzed its utility. Compared to state-of-the-art solutions in the central model of DP, FastLloyd results in higher utility across a wide range of real datasets and scales effectively to larger dimensions and number of clusters. FastLloyd also achieves four orders of magnitude faster runtime than the state-of-the-art in secure federated $k$-means across a variety of problem sizes. In summary, we provide an efficient, private, and accurate solution to the horizontally federated $k$-means problem.
Acknowledgements.
We gratefully acknowledge the support of the National Research Council Canada, the University of Waterloo, the Government of Ontario, and the NSERC Postgraduate Scholarship-Doctoral program for funding this research.
Appendix A Proof of Quantized DP Noise

### v2

@@ 6.1.3 Datasets
Iris [23] 150 4 3
LSun [65] 400 2 3
S1 [24] 5000 2 15
House [25] 1837 3 3
Adult [7] 48842 6 3
Wine [1] 178 13 3
Breast [67] 699 9 2
Yeast [51] 1484 8 10
MNIST [41] 60000 784 10
Birch2 [72] 25000 2 100
G2 [46] 2048 2-1024 2
Synth 10K 2-512 2-32
Synth-K 10K 2 2-128
@@
(a) Varying the number of dimensions using G2 ($k=2$, log scale).
(b) Varying the number of clusters using Synth-K ($d=2$).
Figure 3: Scaling of AUC over number of dimensions and clusters.
@@
(c) Improvement
Figure 4: AUC comparison of FastLloyd vs. SuLloyd on the Synth datasets.
6.2 Utility Evaluation
6.2.1 Metrics
We use the Normalized Intra-cluster Variance (NICV) from prior work [64] as our primary utility metric. NICV normalizes the $k$-means objective function (Eqn 1) by dividing it by the dataset size:
$NICV(O)=\frac{1}{N}\cdot\sum_{j=1}^{k}\sum_{x_{l}\in O_{j}}||x_{l}-\mu_{j}||^{2}$
To facilitate systematic comparison across methods for the synthetic datasets, we summarize each method’s performance over different privacy budgets using the area under the curve (AUC) of NICV values against $\epsilon\in[0.1,0.25,0.5,0.75,1.0]$, computed via the trapezoidal rule:
$AUC=\sum_{i=1}^{n-1}\frac{NICV_{i}+NICV_{i+1}}{2}\cdot(\epsilon_{i+1}-\epsilon_{i})$ (26)
6.2.2 Real World Datasets
Figure 2 shows the NICV against various privacy budgets on ten real-world datasets to answer Q1. The shading shows the 95% confidence interval of the mean over the 100 experiments. We observe that FastLloyd and GLloyd consistently outperform SuLloyd across all high-dimensional datasets (e.g., Wine, Breast, Yeast, MNIST). This aligns with our theoretical expectations as the Gaussian mechanism composes much more favourably in high-dimensional spaces than the Laplace mechanism. Additionally, FastLloyd outperforms GLloyd in all cases, especially in datasets with low dimensions and a high number of clusters (e.g., Birch2, S1), where GLloyd does not outperform SuLloyd. This highlights the effect of imposing a radius constraint on the protocol, which, as discussed in Section 5.1, is most restrictive in low-dimensional, high-cluster settings, further reducing the noise added to the centroids. While all differentially private methods eventually approach the non-private baseline as privacy budget increases, FastLloyd achieves this convergence at substantially lower privacy budgets, establishing it as the superior approach for DP $k$-means clustering across all datasets.
6.2.3 Synthetic Datasets
To answer Q2, we follow an approach similar to Su et al. [64] and evaluate FastLloyd and SuLloyd on our synthetic datasets varying both $k$ and $d$. Figure 4 shows heatmaps of the AUC values for SuLloyd and FastLloyd on the synthetic datasets. To clarify the difference between the protocols, Figure 4(c) shows the percentage of reduction in AUC of FastLloyd over that of SuLloyd. From the heatmaps, we observe that FastLloyd outperforms SuLloyd across all datasets, with the improvement being more pronounced in higher dimensions and in higher $k$.
To investigate higher limits, we further extend the evaluation to the G2 [46] datasets with $d$ up to 1024, and Synth-K datasets with $k$ up to 128. Figures 3(a) and 3(b) present a comparative analysis of AUC performance across Lloyd, FastLloyd, GLloyd, and SuLloyd protocols, evaluated on G2 and Synth-K datasets. The AUC values are normalized relative to SuLloyd as the baseline. Figure 3(a) examines dimensionality scaling by varying $d$ up to 1024 while maintaining $k=2$, whereas Figure 3(b) shows cluster scaling by varying $k$ up to 128 while fixing $d=2$. In Figure 3(a), we observe that for low dimensions, all protocols perform very closely to the baseline (Lloyd), but as the dimension increases, the task becomes much more challenging for the private protocols, with FastLloyd outperforming all others. In Figure 3(b), we observe that FastLloyd outperforms all other protocols across all cluster counts, with the improvement being more pronounced at higher cluster counts. This is to be expected as the radius constraint becomes more restrictive with higher cluster counts, leading to a more significant reduction in noise added to the centroids.
Since Su et al. found SuLloyd outperformed all other approaches for $d{}>3$ [64], and we outperform SuLloyd, we claim that FastLloyd is state-of-the-art in DP $k$-means for $d{}>3$. FastLloyd also fixes the scalability issue that Su et al.’s work [64] observed in high $k$ and enables us to scale to much larger dimensions.
@@ 6.2.3 Synthetic Datasets
$N$ $k$ $d$ FastLloyd MohLloyd [50] Speedup FastLloyd MohLloyd [50] Reduction
TimeSynth 10K 2 2 $3.03\pm 0.02$ 10500 3465×3465$\times$3465 × 192 2.557e8 $1.33e6\times$
2 5 $3.24\pm 0.01$ 384 - -
5 2 $4.32\pm 0.02$ 34050 7882×7882$\times$7882 × 480 9.746e8 $2.03e6\times$
5 5 $4.95\pm 0.04$ 960 - -
TimeSynth 100K 2 2 $12.57\pm 0.02$ 105120 8363×8363$\times$8363 × 192 2.467e9 $1.28e7\times$
2 5 $13.17\pm 0.03$ 384 - -
5 2 $22.05\pm 0.03$ 347250 15748×15748$\times$15748 × 480 9.535e9 $1.99e7\times$
5 5 $22.56\pm 0.02$ 960 - -
LSun 400 3 2 $2.6\pm 0.04$ 1481 570×570$\times$570 × 288 - -
S1 5K 15 2 $5.81\pm 0.08$ 49087 8449×8449$\times$8449 × 1440 - -
@@
Table 3: Overhead comparison per iteration against MohLloyd [50] for two clients ($100$ runs with mean and $95\%$ confidence reported for FastLloyd)
6.3 Runtime Evaluation
@@
To answer questions Q3 and Q4, we compare FastLloyd with the MohLloyd protocol of Mohassel et al. [50] on a variety of common datasets in terms of runtime and communication size (per iteration). Table 3 shows the summary of the evaluation with two clients (the setting considered in Mohassel et al.’s work [50]). Values for MohLloyd are taken as reported in the paper by Mohassel et al. [50], noting that the setup is similar to the one used in our evaluation, and the gap in performance is more than what could be accounted for by different setups. In terms of runtime, MohLloyd executes in the order of minutes, while FastLloyd executes in milliseconds, offering five orders of magnitude speedup. In terms of communication size, MohLloyd requires communicating gigabytes of data per iteration, while FastLloyd needs a fraction of a kilobyte, offering up to seven orders of magnitude reduction in size. In terms of communication rounds, MohLloyd requires $\Theta(\lceil\log{k{}}\rceil)$ communication rounds per iteration [31], while FastLloyd only requires one. Since MohLloyd was found to be the state-of-the-art in secure federated $k$-means [31], we conclude that we advance the state-of-the-art while offering output privacy (which would only further slow down Mohassel et al. [50]).
7 Conclusion
In this work, we designed FastLloyd, a new private protocol for federated $k$-means. We have shown that FastLloyd is secure in the computational model of DP and analyzed its utility. Compared to state-of-the-art solutions in the central model of DP, FastLloyd results in higher utility across a wide range of real datasets and scales effectively to larger dimensions and number of clusters. FastLloyd also achieves five orders of magnitude faster runtime than the state-of-the-art in secure federated $k$-means across a variety of problem sizes. In summary, we provide an efficient, private, and accurate solution to the horizontally federated $k$-means problem.
Acknowledgments
We gratefully acknowledge the support of NSERC for grants RGPIN-2023-03244, IRC-537591, the Government of Ontario and the Royal Bank of Canada for funding this research.
Appendix A Proofs
A.1 Proof of Theorem 5.1
See 5.1
@@ Appendix A Proofs
w.l.o.g assume that the datasets differ by a single point $x^{\prime}$ so that $D^{\prime}=D\cup\{x^{\prime}\}$. If $x^{\prime}$ is not within $\eta$ of $\tilde{\mu}_{j}^{(t-1)}$, it will not be assigned to $O_{j}$ and thus $R_{j}(D)=R_{j}(D^{\prime})$. Therefore, we only need to consider the case where $x^{\prime}$ is at most $\eta$ from $\tilde{\mu}_{j}^{(t-1)}$. By definition, we have:
$\Delta^{R}$ $=$ $\max\limits_{D,D^{\prime}\in\mathcal{D}}||S_{j}^{(t)}-C_{j}^{(t)}\tilde{\mu}_{j}^{(t-1)}-(S_{j}^{\prime(t)}-C_{{}^{\prime}}{j}^{(t)}\tilde{\mu}_{j}^{(t-1)})||_{2}$
$=$ $\max\limits_{D,D^{\prime}\in\mathcal{D}}||S_{j}^{(t)}-C_{j}^{(t)}\tilde{\mu}_{j}^{(t-1)}-S_{j}^{(t)}-x^{\prime}+(C_{j}^{(t)}+1)\tilde{\mu}_{j}^{(t-1)}||_{2}$
$=$ $\max\limits_{D,D^{\prime}\in\mathcal{D}}||\tilde{\mu}_{j}^{(t-1)}-x^{\prime}||_{2}$
$\leq$ $\eta$
where the last line follows by (Eqn 4). ∎
A.2 Proof of Theorem 5.2
We first prove the following helpful lemma.
Lemma A.1.
Our noise mechanism (defined in Eqn 8) of $f(D)+\gamma$ where $\gamma\sim\mathcal{N}(0,\sigma^{2}T(\Delta^{(f)})^{2})$ applied over $T$ adaptive iterations of $f$ is $(\frac{1}{\sigma})$-GDP, where $\Delta^{(f)}$ is the sensitivity of $f$.
Proof.
After choosing $\sigma$, by Theorem 2.1 from Dong et al. [15, Theorem 1], we have that each application of a Gaussian mechanism with noise multiplier $\sigma$ is $\frac{\Delta^{(f)}}{\sigma}-GDP$. We apply the Gaussian mechanism adaptively over $T$ iterations. Which by Theorem 2.2 from Dong et al. [15, Corollary 2], gives us $\frac{\Delta^{(f)}\sqrt{T}}{\sigma}$-GDP. Thus, by multiplying $\sigma$ by $\sqrt{T}\Delta^{(f)}$, we get $\frac{1}{\sigma}$-GDP. ∎
We now restate and prove Theorem 5.2. See 5.2
Proof.
To prove the algorithm satisfies Definition 3.1, we need to consider the view from each party. We begin with the view of the server. First, let us assume that each client samples the random mask $r_{i}$ from a random oracle over $\mathbf{R}$. Under this assumption, the $Enc(\cdot)$ function satisfies information-theoretic security as it is a one-time pad using $r_{i}$ as the pad. In practice, we implement the sampling of $r_{i}$ using a PRNG and thus reduce from information-theoretic to computational security with a negligible term $negl(\lambda)$. Therefore, the view of the server satisfies $\epsilon(\lambda)$-IND-CDP-MPC with $\epsilon=0$.
Regarding the clients, a client’s view consists of their own dataset $D_{A}$, and the output from each iteration of the protocol. We note that the initialization is data-independent and thus is indistinguishable between neighbouring datasets. Each iteration consists of the assignment and updating of the cluster centroids. The assignment step uses the published centroids from the previous iteration (post-processing) to divide the dataset into clusters. We can then apply parallel composition over each of the clusters. Thus, we can focus on the privacy cost of a single cluster for the remainder of the proof.
First, we set $\sigma$ using Theorem 2.3 from Dong et al. [15, Corollary 1]. We choose the minimum $\sigma$ such that $(\epsilon,\delta)$-DP iff $\frac{1}{\sigma}$-GDP by Theorem 2.3. This is equivalent to finding the noise multiplier for a sensitivity one, single application, of the Gaussian mechanism with noise multiplier $\sigma$. To solve this minimization, we use Algorithm 1 from Balle and Wang [6].
The updating of the cluster centroid applies the Gaussian mechanism twice, once to the relative sums and once to the counts. We split the overall $\sigma$ into $\sigma^{R}$ and $\sigma^{C}$ following Eqn 7. We begin with the analysis of the relative sum. In Theorem 5.1, we show that the sensitivity of the relative sum is $\eta$. Applying Lemma A.1 we get that the relative sum over $T$ iterations is $\frac{1}{\sigma^{R}}$-GDP. The sensitivity of the count is $1$, and thus the count is $\frac{1}{\sigma^{C}}$-GDP by similar analysis. Finally, applying Theorem 2.2. We get that
$\sqrt{\left(\frac{1}{\sigma^{R}}\right)^{2}+\left(\frac{1}{\sigma^{C}}\right)^{2}}=\frac{1}{\sigma}$ (27)
and thus the we get $\frac{1}{\sigma}$-GDP over the entire protocol which implies $(\epsilon,\delta)$-DP (because of how we chose $\sigma$). Applying parallel composition over all clusters, the client’s view is $(\epsilon,\delta)$-IND-CDP-MPC, with no computational assumption (as we use the information-theoretic DP properties of differential privacy). Taking the worst case over the clients and the server, the result follows. ∎
A.3 Proof of Quantized DP Noise